# Université de Genève, CH-1211 Genève, Switzerland

### v1

@@ 2.2 The case of $\text{SL}(2,{\mathbb{C}})$ gauge group
In this work, we take $G_{{\mathbb{C}}}=\text{SL}(2,{\mathbb{C}})$. Therefore, we have simply $r=1$, and
@@
is the classical $A$-polynomial of the knot Cooper:1994aa . The operators $\hat{P},\hat{X}$ act on $u$-states as
$\hat{P}|u\rangle=|u+{\rm i}{\mathsf{b}}\rangle\,,\quad\hat{X}|u\rangle=X|u\rangle\,.$ (2.19)
Equivalently, $\hat{P}$ is the operator that shifts $X$ into $qX$, and $\hat{X}$ is the ordinary multiplication. Building on previous results hikami01 ; hikami07 ; dglz , the exact, non-perturbative partition function $\chi_{{\cal K}}(u;\hbar)$ has been constructed from an ideal triangulation of $M$ and identified with a finite-dimensional integral whose integrand is a product of Faddeev’s quantum dilogarithm functions faddeev , known as the state-integral invariant, or Andersen–Kashaev invariant ak ; ak2 ; dimofte-rev . This is a holomorphic function of =\$\0\hbar\in{\mathbb{C}}^{\prime}={\mathbb{C}}\backslash(-\infty,0]$ and $u\in{\mathbb{C}}$. It can be explicitly verified amalusa that it obeys the $q$-difference equation encoded in the quantized $A$-polynomial of the knot, that is,
$\hat{A}_{{\cal K}}(\hat{P},\hat{X},q)\chi_{{\cal K}}(u;\hbar)=0\,.$ (2.20)
@@
$\hbar_{D}={4\pi^{2}\over\hbar}=\frac{2\pi}{\tau}\,,\quad q_{D}={\rm e}^{{\rm i}\hbar_{D}}={\rm e}^{\frac{2\pi{\rm i}}{\tau}}=\tilde{q}^{-1},$ (2.21)
@@
$X_{D}={\rm e}^{\frac{2\pi u}{{\mathsf{b}}}}=X^{2\pi\over\hbar}\,.$ (2.22)
@@
and, in particular, it satisfies the $S$-dual quantum $A$-polynomial equation
$\hat{A}_{\cal K}(\hat{P}_{D},\hat{X}_{D},q_{D})\chi_{{\cal K}}(u;\hbar)=0\,,$ (2.24)
@@
Throughout this work, we look at two concrete examples of hyperbolic knots, which allow us to perform explicit and detailed calculations to provide support to and give insight into our claims. These are the figure-eight and the three-twist knots.
@@
The figure-eight knot, also denoted by $\bm{4}_{1}$, is shown on the left side of Fig. 1. Its classical $A$-polynomial can be written in symmetrized form as
$A_{\bm{4}_{1}}(P,X)=P^{-1}+P-(X^{2}+X^{-2}-X-X^{-1}-2)\,.$ (2.25)
@@
$P^{(\rm{geom},\rm{conj})}=\frac{T(X)\pm(1-X^{2})\sqrt{\Delta(X)}}{2X^{2}}\,,$ (2.27)
@@
Building on the results and conventions of ggm2 , the descendant $\hat{A}$-operator is given by
$\hat{A}_{\bf{4}_{1}}(\hat{P},\hat{X},q^{m},q)=\sum_{j=0}^{2}C_{j}(\hat{X},q^{m},q)\hat{P}^{j}\,,$ (2.29)
where $m$ takes values in ${\mathbb{Z}}$ and the coefficient functions $C_{j}(\hat{X},q^{m},q)$ are
$C_{0}(\hat{X},q^{m},q)=$ $q^{2+3m}\hat{X}^{2}(-1+q^{3+m}\hat{X}^{2})\,,$ (2.30a)
$C_{1}(\hat{X},q^{m},q)=$ $-q^{m}(-1+q^{2+m}\hat{X}^{2})(1-q\hat{X}-q^{1+m}\hat{X}^{2}$ (2.30b)
$-q^{3+m}\hat{X}^{2}-q^{3+m}\hat{X}^{3}+q^{4+2m}\hat{X}^{4})\,,$
$C_{2}(\hat{X},q^{m},q)=$ $q^{2}\hat{X}^{2}(-1+q^{1+m}\hat{X}^{2})\,.$ (2.30c)
@@
$\hat{A}_{\bf{4}_{1}}(\hat{P},\hat{X},q)=C_{0}(\hat{X},q)\hat{P}^{-1}+C_{1}(\hat{X},q)+C_{2}(\hat{X},q)\hat{P}\,,$ (2.31)
where the coefficient functions $C_{j}(\hat{X},q)$ are
$C_{0}(\hat{X},q)=$ $q\hat{X}^{2}-1\,,$ (2.32a)
$C_{1}(\hat{X},q)=$ $-(\hat{X}^{2}-1)(\hat{X}^{2}+\hat{X}^{-2}-\hat{X}-\hat{X}^{-1}-q-q^{-1})\,,$ (2.32b)
$C_{2}(\hat{X},q)=$ $q^{-1}\hat{X}^{2}-1\,.$ (2.32c)
@@
The three-twist knot, also denoted by $\bm{5}_{2}$, is shown on the right side of Fig. 1. Its classical $A$-polynomial can be written in symmetrized form as
$A_{\bm{5}_{2}}(P,X)=$ $P^{-1}-(X^{-2}-X^{-1}+2X+2X^{2}-X^{3})$ (2.33)
$-(X^{-1}-2-2X+X^{3}-X^{4})P-X^{2}P^{2}\,.$
@@
$\alpha=\rm{geom},\rm{conj},\rm{\text{self}}\,.$ (2.34)
Again, following ggm2 , the descendant $\hat{A}$-operator is given by
$\hat{A}_{\bf{5}_{2}}(\hat{P},\hat{X},q^{m},q)=\sum_{j=0}^{3}C_{j}(\hat{X},q^{m},q)\hat{P}^{j}\,,$ (2.35)
where $m$ takes values in ${\mathbb{Z}}$ and the coefficient functions $C_{j}(\hat{X},q^{m},q)$ are
$C_{0}(\hat{X},q^{m},q)=$ $-q^{2+m}\hat{X}^{2}(1-q^{2}\hat{X})(1+q^{2}\hat{X})(1-q^{5}\hat{X}^{2})\,,$ (2.36a)
$C_{1}(\hat{X},q^{m},q)=$ $(1-q\hat{X})(1+q\hat{X})(1-q^{5}\hat{X}^{2})(1-q\hat{X}-q\hat{X}^{2}-q^{4}\hat{X}^{2}$ (2.36b)
$+q^{2+m}\hat{X}^{2}+q^{3+m}\hat{X}^{2}+q^{2}\hat{X}^{3}+q^{5}\hat{X}^{3}+q^{5}\hat{X}^{4}+q^{5+m}\hat{X}^{4}-q^{6}\hat{X}^{5})\,,$
$C_{2}(\hat{X},q^{m},q)=$ $q\hat{X}(1-q^{2}\hat{X})(1+q^{2}\hat{X})(1-q\hat{X}^{2})(1-q^{2}\hat{X}-q^{2+m}\hat{X}-q^{2}\hat{X}^{2}-q^{5}\hat{X}^{2}$ (2.36c)
$+q^{4}\hat{X}^{3}+q^{7}\hat{X}^{3}-q^{5+m}\hat{X}^{3}-q^{6+m}\hat{X}^{3}+q^{7}\hat{X}^{4}-q^{9}\hat{X}^{5})\,,$
$C_{3}(\hat{X},q^{m},q)=$ $q^{8+m}\hat{X}^{4}(1-q\hat{X})(1+q\hat{X})(1-q\hat{X}^{2})\,.$ (2.36d)
@@
$\hat{A}_{\bf{5}_{2}}(\hat{P},\hat{X},q)=C_{0}(\hat{X},q)\hat{P}^{-2}+C_{1}(\hat{X},q)\hat{P}^{-1}+C_{2}(\hat{X},q)+C_{3}(\hat{X},q)\hat{P}\,,$ (2.37)
@@
$C_{0}(\hat{X},q)=$ $-\hat{X}^{2}(1-q^{2}\hat{X}^{2})(1-q^{3}\hat{X}^{2})\,,$ (2.38a)
$C_{1}(\hat{X},q)=$ $(1-\hat{X}^{2})(1-q^{3}\hat{X}^{2})(q^{2}\hat{X}^{3}-q^{2}\hat{X}^{2}-q\hat{X}^{5}$ (2.38b)
$+2q\hat{X}^{4}+q^{-1}\hat{X}^{3}+q\hat{X}^{2}-q^{-1}\hat{X}^{2}+\hat{X}^{2}-\hat{X}+1)\,,$
$C_{2}(\hat{X},q)=$ $\hat{X}(1-q^{2}\hat{X}^{2})(1-q^{-1}\hat{X}^{2})(-q^{4}\hat{X}^{5}+q^{4}\hat{X}^{3}$ (2.38c)
$+q^{3}\hat{X}^{4}-q^{3}\hat{X}^{3}-q^{3}\hat{X}^{2}-q^{2}\hat{X}^{3}+q\hat{X}^{3}-2q\hat{X}-\hat{X}^{2}+1)\,,$
$C_{3}(\hat{X},q)=$ $q^{4}\hat{X}^{4}(1-\hat{X}^{2})(1-q^{-1}\hat{X}^{2})\,.$ (2.38d)
@@ 3 The structure of the wave function
In this section, we study the non-perturbative wave function $\chi_{{\cal K}}(u;\hbar)$ introduced in Section 2 by exploiting ideas from topological string theory. In particular, we propose various new conjectures on the structure of this wave function, which are motivated by similar results in the open version of the TS/ST correspondence kpamir ; mz-wv ; mz-wv2 ; szabolcs and by our detailed analysis of the examples presented in Section 4. In the following, we will not explicitly indicate the knot ${\cal K}$ to avoid the cluttering of notation.
@@ 3.1 General conjectures
Our first claim is that the WKB expansion in Eq. (2.17) can be resummed at all orders in $\hbar$ and order by order in $X={\rm e}^{2\pi{\mathsf{b}}u}$. Explicitly, for a fixed choice of $\alpha$, we have that
@@
$a^{(\alpha)}_{\pm n}(q)={\hat{a}_{\pm n}^{(\alpha)}(q)\over n(q^{n}-1)q^{\sigma_{\alpha,\pm n}}}\,,\quad n\in{\mathbb{Z}}_{>0}\,,$ (3.41)
@@
$\hat{a}^{(\rm{geom})}_{-n}(q)=0\,,\quad n\in{\mathbb{Z}}_{>0}\,,$ (3.42)
and $\sigma_{\alpha,m}$ is a function of $m\in{\mathbb{Z}}_{\neq 0}$ which takes values in ${\mathbb{N}}$, depends on the choice of classical branch $\alpha$, and satisfies
$\sigma_{{\rm geom},m}=0\,,\quad\sigma_{\alpha,-|m|}=0\,\quad\forall\alpha\,.$ (3.43)
Note that we use the same notation for the perturbative WKB series in Eq. (2.17) and its resummation in terms of $q$ in Eq. (3.39). However, in the rest of this paper, we will always refer to the resummed version of $\chi^{(\alpha),\rm{WKB}}(u;\hbar)$ as in Eq. (3.39). In fact, there is additional structure in the resummed perturbative wave function. Our second claim is that $\phi^{(\alpha),\rm{WKB}}(X;q)$ has an integrality/multicovering-type structure. Namely, for a fixed choice of $\alpha$, we have that
$\phi^{(\alpha),\rm{WKB}}(X;q)=\sum_{k,s\geq 1}{D^{(\alpha)}_{s}(q^{k})\over k(q^{k}-1)q^{k\sigma_{\alpha,s}}}X^{sk}+{D^{(\alpha)}_{-s}(q^{k})\over k(q^{k}-1)q^{k\sigma_{\alpha,-s}}}X^{-sk}\,,$ (3.44)
where $D^{(\alpha)}_{\pm s}(q)\in{\mathbb{Z}}[q]$. Equivalently, we can write the coefficient functions from Eq. (3.40) as
$a^{(\alpha)}_{\pm n}(q)=\sum_{k|n}{D^{(\alpha)}_{\pm{n\over k}}(q^{k})\over k(q^{k}-1)q^{k\sigma_{\alpha,\pm{n\over k}}}}\,,\quad n\in{\mathbb{Z}}_{>0}\,,$ (3.45)
@@
${\tilde{D}}_{\pm s}(q)={D_{\pm s}(q)\over(q-1)q^{\sigma_{\pm s}}}\,,\quad s\in{\mathbb{Z}}_{>0}\,.$ (3.46)
As a consequence of Eq. (3.44), we can write
$\exp\left(\phi^{\rm WKB}(X;q)\right)=$ $\exp\left(\sum_{k,s\geq 1}{1\over k}{\tilde{D}}_{s}(q^{k})X^{sk}\right)\exp\left(\sum_{k,s\geq 1}{1\over k}{\tilde{D}}_{-s}(q^{k})X^{-sk}\right)$ (3.47)
$=$ ${\rm Exp}_{X,q}\left(\sum_{s\geq 1}{\tilde{D}}_{s}(q)X^{s}\right){\rm Exp}_{X,q}\left(\sum_{s\geq 1}{\tilde{D}}_{-s}(q)X^{-s}\right)\,,$
where ${\rm Exp}_{X,q}$ is the plethystic exponential in the variables $X,q$. See, e.g.lm-pv . It follows that we can extract the functions ${\tilde{D}}_{\pm s}(q)$ by applying the plethystic logarithm in the same two variables, which we denote by ${\rm Log}_{X,q}$, to both sides of Eq. (3.47). Specifically, we find that
$\sum_{s\geq 1}{\tilde{D}}_{\pm s}(q)X^{\pm s}={\rm Log}_{X,q}\left(\exp\left(\sum_{n\geq 1}a_{\pm n}(q)X^{\pm n}\right)\right)=\sum_{k,n\geq 1}{\mu(k)\over k}a_{\pm n}(q^{k})X^{\pm nk}\,,$ (3.48)

### v2

@@ 2.2 The case of $\bm{\mathrm{SL}(2,{\mathbb{C}})}$ gauge group
In this work, we take $G_{{\mathbb{C}}}=\mathrm{SL}(2,{\mathbb{C}})$. Therefore, we have simply $r=1$ and
@@
is the classical $A$-polynomial of the knot [7]. The operators $\hat{P}$, $\hat{X}$ act on $u$-states as
$\hat{P}|u\rangle=|u+{\rm i}{\mathsf{b}}\rangle,\qquad\hat{X}|u\rangle=X|u\rangle.$ (2.20)
Equivalently, $\hat{P}$ is the operator that shifts $X$ into $qX$, and $\hat{X}$ is the ordinary multiplication. Building on previous results [11, 26, 27], the exact, non-perturbative partition function $\chi_{{\cal K}}(u;\hbar)$ has been constructed from an ideal triangulation of $M$ and identified with a finite-dimensional integral whose integrand is a product of Faddeev’s quantum dilogarithm functions [13], known as the state-integral invariant, or Andersen–Kashaev invariant [2, 3, 9]. This is a holomorphic function of $\hbar\in{\mathbb{C}}^{\prime}={\mathbb{C}}\backslash(-\infty,0]$ and $u\in{\mathbb{C}}$. It has been explicitly verified in the examples of the figure-eight and three-twist knots and conjectured in general [4] that it obeys the $q$-difference equation encoded in the quantum $A$-polynomial of the knot, that is,
$\hat{A}_{{\cal K}}\bigl(\hat{P},\hat{X},q\bigr)\chi_{{\cal K}}(u;\hbar)=0.$ (2.21)
@@
$\hbar_{D}={4\pi^{2}\over\hbar}={{2\pi\over\tau}},\qquad q_{D}={\rm e}^{{\rm i}\hbar_{D}}={\rm e}^{{{2\pi{\rm i}\over\tau}}}=\tilde{q}^{-1},$ (2.22)
@@
$X_{D}={\rm e}^{{{2\pi u\over{\mathsf{b}}}}}=X^{2\pi\over\hbar}.$ (2.23)
@@
and, in particular, it satisfies the $S$-dual quantum $A$-polynomial equation [^5] Since the wave function simultaneously obeys distinct $q$- and $q_{D}$-difference equations, the solution space of equations (2.21) and (2.25) may be naturally regarded as a bimodule over the difference operator algebras associated with the quantum $A$-polynomial and its $S$-dual image. We will not, however, pursue this perspective further here.
$\hat{A}_{{\cal K}}\bigl(\hat{P}_{D},\hat{X}_{D},q_{D}\bigr)\chi_{{\cal K}}(u;\hbar)=0,$ (2.25)
@@
Throughout this work, we examine the two simplest examples of hyperbolic knots, which allow us to perform explicit and detailed calculations to support and give insight into our claims. These are the figure-eight and three-twist knots.
@@
The figure-eight knot, denoted as $\bm{4}_{1}$, is shown on the left side of Figure 1. Its classical $A$-polynomial can be written in symmetrized form as
$A_{\bm{4}_{1}}(P,X)=P^{-1}+P-\bigl(X^{2}+X^{-2}-X-X^{-1}-2\bigr).$ (2.26)
@@
$P^{({\rm geom},{\rm conj})}={{T(X)\pm\bigl(1-X^{2}\bigr)\sqrt{\Delta(X)}\over 2X^{2}}},$ (2.28)
@@
Building on the results and conventions of [17], the descendant $\hat{A}-$operator is given by
$\hat{A}_{\bf{4}_{1}}\bigl(\hat{P},\hat{X},q^{m},q\bigr)=\sum_{j=0}^{2}C_{j}\bigl(\hat{X},q^{m},q\bigr)\hat{P}^{j},$ (2.30)
where $m$ takes values in $\mathbb{Z}$} and the coefficient functions $C_{j}\bigl(\hat{X},q^{m},q\bigr)$ are
$C_{0}\bigl(\hat{X},q^{m},q\bigr)=q^{2+3m}\hat{X}^{2}\bigl(-1+q^{3+m}\hat{X}^{2}\bigr),$ (2.31a)
$C_{1}\bigl(\hat{X},q^{m},q\bigr)=-q^{m}\bigl(-1+q^{2+m}\hat{X}^{2}\bigr)\bigl(1-q\hat{X}-q^{1+m}\hat{X}^{2}$
[…]
$C_{2}\bigl(\hat{X},q^{m},q\bigr)=q^{2}\hat{X}^{2}\bigl(-1+q^{1+m}\hat{X}^{2}\bigr).$ (2.31c)
@@
$\hat{A}_{\bf{4}_{1}}\bigl(\hat{P},\hat{X},q\bigr)=C_{0}\bigl(\hat{X},q\bigr)\hat{P}^{-1}+C_{1}\bigl(\hat{X},q\bigr)+C_{2}\bigl(\hat{X},q\bigr)\hat{P}$, (2.32)
where the coefficient functions $C_{j}\bigl(\hat{X},q\bigr)$ are
$C_{0}\bigl(\hat{X},q\bigr)=q\hat{X}^{2}-1,$ (2.33a)
$C_{1}\bigl(\hat{X},q\bigr)=-\bigl(\hat{X}^{2}-1\bigr)\bigl(\hat{X}^{2}+\hat{X}^{-2}-\hat{X}-\hat{X}^{-1}-q-q^{-1}\bigr),$ (2.33b)
$C_{2}\bigl(\hat{X},q\bigr)=q^{-1}\hat{X}^{2}-1.$ (2.33c)
@@
The three-twist knot, denoted as $\bm{5}_{2}$, is shown on the right side of Figure 1. Its classical $A$-polynomial can be written in symmetrized form as
$A_{\bm{5}_{2}}(P,X)=P^{-1}-\bigl(X^{-2}-X^{-1}+2X+2X^{2}-X^{3}\bigr)$
[…]
@@
$\alpha={\rm geom},{\rm conj},{\rm self}.$ (2.35)
Again, following [17], the descendant $\hat{A}-$operator is given by
$\hat{A}_{\bf{5}_{2}}\bigl(\hat{P},\hat{X},q^{m},q\bigr)=\sum_{j=0}^{3}C_{j}\bigl(\hat{X},q^{m},q\bigr)\hat{P}^{j},$ (2.36)
where $m$ takes values in $\mathbb{Z}$} and the coefficient functions $C_{j}\bigl(\hat{X},q^{m},q\bigr)$ are
$C_{0}\bigl(\hat{X},q^{m},q\bigr)={}$ $-q^{2+m}\hat{X}^{2}\bigl(1-q^{2}\hat{X}\bigr)\bigl(1+q^{2}\hat{X}\bigr)\bigl(1-q^{5}\hat{X}^{2}\bigr),$ (2.37a)
$C_{1}\bigl(\hat{X},q^{m},q\bigr)={}$ $\bigl(1-q\hat{X}\bigr)\bigl(1+q\hat{X}\bigr)\bigl(1-q^{5}\hat{X}^{2}\bigr)\bigl(1-q\hat{X}-q\hat{X}^{2}-q^{4}\hat{X}^{2}+q^{2+m}\hat{X}^{2}$
$+q^{3+m}\hat{X}^{2}+q^{2}\hat{X}^{3}+q^{5}\hat{X}^{3}+q^{5}\hat{X}^{4}+q^{5+m}\hat{X}^{4}-q^{6}\hat{X}^{5}\bigr),$ (2.37b)
$C_{2}\bigl(\hat{X},q^{m},q\bigr)={}$ $q\hat{X}\bigl(1-q^{2}\hat{X}\bigr)\bigl(1+q^{2}\hat{X}\bigr)\bigl(1-q\hat{X}^{2}\bigr)\bigl(1-q^{2}\hat{X}-q^{2+m}\hat{X}-q^{2}\hat{X}^{2}-q^{5}\hat{X}^{2}$
$+q^{4}\hat{X}^{3}+q^{7}\hat{X}^{3}-q^{5+m}\hat{X}^{3}-q^{6+m}\hat{X}^{3}+q^{7}\hat{X}^{4}-q^{9}\hat{X}^{5}\bigr),$ (2.37c)
$C_{3}\bigl(\hat{X},q^{m},q\bigr)={}$ $q^{8+m}\hat{X}^{4}\bigl(1-q\hat{X}\bigr)\bigl(1+q\hat{X}\bigr)\bigl(1-q\hat{X}^{2}\bigr).$ (2.37d)
@@
$\hat{A}_{\bf{5}_{2}}\bigl(\hat{P},\hat{X},q\bigr)=C_{0}\bigl(\hat{X},q\bigr)\hat{P}^{-2}+C_{1}\bigl(\hat{X},q\bigr)\hat{P}^{-1}+C_{2}\bigl(\hat{X},q\bigr)+C_{3}\bigl(\hat{X},q\bigr)\hat{P}$, (2.38)
@@
$C_{0}\bigl(\hat{X},q\bigr)={}$ $-\hat{X}^{2}\bigl(1-q^{2}\hat{X}^{2}\bigr)\bigl(1-q^{3}\hat{X}^{2}\bigr),$ (2.39a)
$C_{1}\bigl(\hat{X},q\bigr)={}$ $\bigl(1-\hat{X}^{2}\bigr)\bigl(1-q^{3}\hat{X}^{2}\bigr)\bigl(q^{2}\hat{X}^{3}-q^{2}\hat{X}^{2}-q\hat{X}^{5}$
$+2q\hat{X}^{4}+q^{-1}\hat{X}^{3}+q\hat{X}^{2}-q^{-1}\hat{X}^{2}+\hat{X}^{2}-\hat{X}+1\bigr),$ (2.39b)
$C_{2}\bigl(\hat{X},q\bigr)={}$ $\hat{X}\bigl(1-q^{2}\hat{X}^{2}\bigr)\bigl(1-q^{-1}\hat{X}^{2}\bigr)\bigl(-q^{4}\hat{X}^{5}+q^{4}\hat{X}^{3}$
$+q^{3}\hat{X}^{4}-q^{3}\hat{X}^{3}-q^{3}\hat{X}^{2}-q^{2}\hat{X}^{3}+q\hat{X}^{3}-2q\hat{X}-\hat{X}^{2}+1\bigr),$ (2.39c)
$C_{3}\bigl(\hat{X},q\bigr)={}$ $q^{4}\hat{X}^{4}\bigl(1-\hat{X}^{2}\bigr)\bigl(1-q^{-1}\hat{X}^{2}\bigr).$ (2.39d)
@@ 3 The structure of the wave function
In this section, we study the non-perturbative wave function $\chi_{{\cal K}}(u;\hbar)$ introduced in Section 2 by exploiting ideas from topological string theory. In particular, we propose new conjectures on the structure of this wave function, which are motivated by similar results in the open version of the TS/ST correspondence [31, 38, 39, 44] and by the evidence presented in Section 4. In the following, we will not explicitly indicate the knot $\cal K$} to avoid the cluttering of notation.
@@ 3.1 General conjectures
Our first claim is that the WKB expansion in equation (2.18) can be resummed at all orders in $\hbar$ and order by order in $X={\rm e}^{2\pi{\mathsf{b}}u}$. Explicitly, for a fixed choice of $\alpha$, we have that
@@
$a^{(\alpha)}_{\pm n}(q)={\hat{a}_{\pm n}^{(\alpha)}(q)\over n(q^{n}-1)q^{\sigma_{\alpha,\pm n}}},\qquad n\in{\mathbb{Z}}_{>0},$ (3.3)
@@
$\hat{a}^{({\rm geom})}_{-n}(q)=0,\qquad n\in{\mathbb{Z}}_{>0},$ (3.4)
and $\sigma_{\alpha,m}$ is a function of $m\in{\mathbb{Z}}_{\neq 0}$ that takes values in $\mathbb{N}$}, depends on the choice of classical branch $\alpha$, and satisfies
$\sigma_{{\rm geom},m}=0,\qquad\sigma_{\alpha,-|m|}=0\qquad\forall\alpha.$ (3.5)
In both examples of the figure-eight and three-twist knots studied in Section 4, we have precise expressions for $\sigma_{\alpha,m}$ for all choices of $\alpha\neq{\rm geom}$, which are given in equations (4.7), (4.27), and (4.28). Moreover, our conjectural statements above simplify notably in the particular case of the geometric branch, whose function $\phi^{({\rm geom}),{\rm WKB}}(X;q)$ has a series expansion including only the positive powers of $X$ and whose coefficients $a^{({\rm geom})}_{n}(q)$ have a trivial factor of $q^{\sigma_{{\rm geom},n}}=1$ in the denominator.
Note that we use the same notation for the perturbative WKB series in equation (2.18) and its resummation in terms of $q$ in equation (3.1). However, in the rest of this paper, we will always refer to the resummed version of $\chi^{(\alpha),{\rm WKB}}(u;\hbar)$ as in equation (3.1). In fact, there is additional structure in the resummed perturbative wave function. Our second claim is that $\phi^{(\alpha),{\rm WKB}}(X;q)$ has an integrality/multicovering-type property. Namely, for a fixed choice of $\alpha$, we have that
$\phi^{(\alpha),{\rm WKB}}(X;q)=\sum_{k,s\geq 1}{D^{(\alpha)}_{s}\bigl(q^{k}\bigr)\over k\bigl(q^{k}-1\bigr)q^{k\sigma_{\alpha,s}}}X^{sk}+{D^{(\alpha)}_{-s}\bigl(q^{k}\bigr)\over k\bigl(q^{k}-1\bigr)q^{k\sigma_{\alpha,-s}}}X^{-sk},$ (3.6)
where $D^{(\alpha)}_{\pm s}(q)\in{\mathbb{Z}}[q]$. Equivalently, we can write the coefficient functions from equation (3.2) as
$a^{(\alpha)}_{\pm n}(q)=\sum_{k|n}{D^{(\alpha)}_{\pm{n\over k}}\bigl(q^{k}\bigr)\over k\bigl(q^{k}-1\bigr)q^{k\sigma_{\alpha,\pm{n\over k}}}},\qquad n\in{\mathbb{Z}}_{>0},$ (3.7)
@@
${\tilde{D}}_{\pm s}(q)={D_{\pm s}(q)\over(q-1)q^{\sigma_{\pm s}}},\qquad s\in{\mathbb{Z}}_{>0}.$ (3.8)
Using equations (3.6) and (3.8), we can write
$\exp\bigl(\phi^{\rm WKB}(X;q)\bigr)$ $=\exp\Bigg(\sum_{k,s\geq 1}{1\over k}{\tilde{D}}_{s}\bigl(q^{k}\bigr)X^{sk}\Bigg)\exp\Bigg(\sum_{k,s\geq 1}{1\over k}{\tilde{D}}_{-s}\bigl(q^{k}\bigr)X^{-sk}\Bigg)$ (3.9)
$=\operatorname{Exp}_{X,q}\Bigg(\sum_{s\geq 1}{\tilde{D}}_{s}(q)X^{s}\Bigg)\operatorname{Exp}_{X,q}\Bigg(\sum_{s\geq 1}{\tilde{D}}_{-s}(q)X^{-s}\Bigg),$
where $\operatorname{Exp}_{X,q}$ denotes the plethystic exponential in the variables $X$, $q$. See, e.g., [35]. It follows that we can extract the functions ${\tilde{D}}_{\pm s}(q)$ by applying the plethystic logarithm in the same two variables, which we denote by ${\rm Log}_{X,q}$, to both sides of equation (3.9). Specifically, we find that
$\sum_{s\geq 1}{\tilde{D}}_{\pm s}(q)X^{\pm s}={\rm Log}_{X,q}\Bigg(\exp\Bigg(\sum_{n\geq 1}a_{\pm n}(q)X^{\pm n}\Bigg)\Bigg)=\sum_{k,n\geq 1}{\mu(k)\over k}a_{\pm n}\bigl(q^{k}\bigr)X^{\pm nk},$ (3.10)